OpenStudy (anonymous):

Carmen buys 153 shares of Cawh Consolidated Banks, each of which pays a constant yearly dividend of $7.14. After six years, how much has Carmen received in dividends?

OpenStudy (mathmale):

Depends upon whether or not this is compound interest. If it's not compound interest, and if each share pays a dividend of $7.14 every year for 6 years, then the total of dividends is \[\frac{ 153 shares }{ 1 }*\frac{ $7.14 }{ 1share*1year }*\frac{ 6 years }{ 1 }\]

OpenStudy (mathmale):

You may cancel shares/share and years/year to obtain a result which is in DOLLARS only.